40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 4 10:30:37 2016, ericq, Update, General, Preliminary vent plans 
    Reply  Tue Jul 5 17:14:02 2016, ericq, Update, General, Vent progress noOsemSide.JPGOSEMside.JPG
       Reply  Tue Jul 5 19:10:41 2016, rana, Update, General, Vent progress jumpEvent20151105.png
          Reply  Tue Jul 5 21:51:35 2016, ericq, Update, General, Vent progress 
             Reply  Tue Jul 5 23:38:42 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation image.jpeg
                Reply  Wed Jul 6 22:58:01 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                   Reply  Thu Jul 7 00:25:07 2016, ericq, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                      Reply  Thu Jul 7 10:49:03 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation UL_glitchMon_Striptool.png
                         Reply  Fri Jul 8 11:35:45 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation coilGlitchMon.png
                            Reply  Fri Jul 8 11:48:09 2016, Koji, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
                               Reply  Fri Jul 8 14:03:26 2016, ericq, Update, General, ETMX removed from suspension zoom.jpg
                                  Reply  Fri Jul 8 16:33:26 2016, ericq, Update, General, ETMX removed from suspension 
                                     Reply  Fri Jul 8 20:44:09 2016, Koji, Update, General, ETMX removed from suspension 6x
                                        Reply  Sun Jul 10 15:38:01 2016, ericq, Update, General, ETMX in guiderod/standoff gluing fixture 6x
                                           Reply  Tue Jul 12 16:08:00 2016, ericq, Update, General, ETMX ruby guided gluing 
                                              Reply  Wed Jul 13 21:13:51 2016, Koji, Update, General, ETMX guide rod gluing / ETMY Magnet gluing 13x
                                                 Reply  Thu Jul 14 22:29:57 2016, ericq , Update, General, ETMX guide rod gluing / ETMY Magnet gluing 
                                                    Reply  Fri Jul 15 16:23:51 2016, ericq , Update, General, Gluing setbacks IMG_2768.JPGIMG_2771.JPG
                                                       Reply  Tue Jul 19 16:39:29 2016, Lydia, Update, General, ETMX magnet gluing/guiderod excess glue removed P7190201.JPGP7190203.JPG
                                                          Reply  Wed Jul 20 13:58:55 2016, Steve, Update, SUS, SUS damping out of vac chamber ETMXrack.jpgfromETMX-satBox.jpgsusDampingSatboxCab.jpgsusDampingSatboxCabl.jpg
                                                          Reply  Thu Jul 21 17:01:59 2016, ericq, Update, General, ETMX magnet gluing/guiderod excess glue removed 
                                                             Reply  Fri Jul 22 23:03:23 2016, gautam, Update, General, ETMX suspension - progress IMG_2828.JPGIMG_2826.JPGIMG_2821.JPGIMG_2822.JPGIMG_2823.JPG
                                                                Reply  Sun Jul 24 03:43:56 2016, ericq, Update, General, Magnet positioning 
                                                                Reply  Mon Jul 25 13:01:49 2016, ericq, Update, General, ETMX suspension - progress IMG_2829.JPGIMG_2830.JPGIMG_2831.JPGIMG_2832.JPGIMG_2833.JPG
                                                                   Reply  Mon Jul 25 13:47:19 2016, rana, Update, General, ETMX suspension - progress 
                                                                   Reply  Mon Jul 25 17:01:54 2016, Steve, Update, SUS, ETMX suspension - progress ITMYside2011.JPG
                                                                   Reply  Mon Jul 25 20:37:46 2016, gautam, Update, General, ETMX suspension - Side OSEM test osem_test_2016_07_25.png
                                                                      Reply  Tue Jul 26 18:07:57 2016, gautam, Update, General, ETMX suspension - Further OSEM Tests OSEM_spectra.pdf
                                                                         Reply  Thu Jul 28 03:22:28 2016, ericq, Update, General, ETMX magnets re-reglued 
                                                                            Reply  Thu Jul 28 13:49:44 2016, gautam, Update, General, ETMX magnets re-reglued ETMX_sideMagnet.JPGETMX_sideMagnet_zoomed.JPG
                                                                               Reply  Thu Jul 28 17:48:17 2016, gautam, Update, General, ETMY LR magnet broke off 
                                                                                  Reply  Thu Jul 28 22:43:03 2016, rana, Update, General, ETMY UL magnet broke off 
                                                                                  Reply  Wed Aug 3 10:52:03 2016, Steve, Update, General, ETMY soaking dish SD1.jpgSD2.jpgSD3.jpg
                                                                               Reply  Fri Jul 29 03:32:30 2016, ericq, Update, General, ETMX suspended once more 
                                                                                  Reply  Fri Jul 29 13:17:34 2016, Koji, Update, General, Oven 
                                              Reply  Thu Jul 14 01:00:33 2016, Koji, Update, General, ETMX ruby guided gluing ruby.jpg
                                  Reply  Thu Jul 28 16:43:01 2016, Lydia, Update, General, ETMX aluminium standoff groove condition wire_damage_4.jpgwire_damage_zoom.jpgno_wire.jpgcomparison_2.jpg
                            Reply  Fri Jul 8 21:22:38 2016, gautam, Update, General, Vent progress - ETMX SUS Coil driver electronics investigation 
             Reply  Wed Jul 6 09:34:26 2016, Steve, Update, General, PRM sensing voltage LR PRM_LR.png
             Reply  Wed Jul 6 19:36:45 2016, Koji, Update, General, SUS Vmon HPF.png160706_ITMX_VMON2.pdf160706_ITMX_VMON1.pdf
                Reply  Thu Jul 7 15:23:39 2016, ericq, Update, General, SUS Vmon coil_drive.png
                   Reply  Thu Jul 7 16:05:55 2016, Koji, Update, General, SUS Vmon 
       Reply  Thu Jul 7 17:12:50 2016, Johannes, Update, General, Vent progress IMG_2621.JPGIMG_2623.JPGIMG_2628.JPGIMG_2595.JPG
          Reply  Tue Jul 12 23:51:16 2016, Johannes, Update, General, Vent progress - ETMY inspection IMG_2679.JPGIMG_2673.JPG
             Reply  Wed Jul 13 00:38:25 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
                Reply  Tue Jul 19 13:21:42 2016, Johannes, Update, General, Vent progress - ETMY attempted repositioning 
       Reply  Sun Jul 10 17:33:00 2016, ericq, Update, General, Vent progress 
          Reply  Mon Jul 11 15:13:22 2016, gautam, Update, General, Vent progress: in-date First Contact procured image.jpeg
Message ID: 12327     Entry time: Fri Jul 22 23:03:23 2016     In reply to: 12322     Reply to this: 12328   12331
Author: gautam 
Type: Update 
Category: General 
Subject: ETMX suspension - progress 

[ericq, gautam]

Summary:

Today, we attempted to progress as far as we could towards getting the mirror suspended and gluing the second wire standoff. We think we have a workable setup now. At this stage, the suspension wire has been looped around the magnet, the second wire standoff has been inserted, coarse pitch balancing has been done, and we have verified that side OSEM/magnet positioning is tenable. Details below.

Details:

  1. First we verified that the epoxy on the side magnet re-glued yesterday had dried (verified using control setup of epoxy in aluminum foil + copper wire - we didn't perform any further tests like pulling the magnet off the tabletop as we were satisfied)
  2. We placed the optic inside the suspension cage, resting on the 4 lower earthquake stops.
  3. We looped the suspension wire around the optic. This is a somewhat challenging procedure. After consulting the documentation, we decided to follow the given advice and loop the wire around from the bottom of the optic, one side at a time. It is tricky to thread the wire between the two lower earthquake stops and get it up around the side. The side magnets were an unexpected ally in this effort as they served as some sort of intermediate checkpoint from which we could pull the wire further up. We then lightly clamped it to the winches mounted atop the suspension cage.
  4. After verifying that we had routed the wire correctly through the various stages (primary and secondary suspension points at the top of the cage), we placed the wire under very slight tension by had, and then tightly clamped the wires in the winches (we then cut off the excess length).
  5. In this state, we proceeded to install the second wire standoff (having verified that the wire was indeed sitting in the groove on the other side). 
  6. We then proceeded to raise the optic to the desired height (center of optic to 5.5 inches above the table top) with the help of the microscope and the lines on the barrel side. 
  7. Next, we attempted to freely suspend the optic (i.e. no contact with the viton tips). We were initially unsuccessful but Eric did some fine adjustment of the (unglued) standoff to achieve a stable configuration. However, the wire is now really close to the magnet - although it is not clear to me if it is touching the magnet as we initially suspected - see Attachment #1, it may be that if the wire is touching something, it is the dumbbell and not the magnet itself. While this is clearly not ideal, we think that this setup is workable as is. If after doing the pitch balancing, if the deviation of the wire becomes much more pronounced, we may have to re-glue the side magnet. In any case, both the horizontal scribed lines are now 5.5 inches above the table top.
  8. We then brought over the OSEMs from the ETMX vacuum chamber to the cleanroom. As a first check, we wanted to ensure that one of the side magnets could accommodate an OSEM (because both side magnets have been re-attached after the optic was removed from the old suspension). Attachment #2 suggests that this is possible, even though the relative positions of the side magnet and the shadow sensor may be sub-optimal. We will only really know after hooking up the electronics.

Remarks:

  1. We found that after a few hours, there was some sag introduced in the wire, presumably it stretched into an equilibrium position under the weight of the optic. We will re-check the heights tomorrow while conducting further tests.

Immediate to-do:

  1. Insert all OSEMS. Ensure that the magnet positions relative to the coil are compatible.
  2. Enable damping loops. We have a cable coming from the IFO area into the cleanroom through a hole-in-the-wall. We are missing a DB25 gender changer at the moment. 
  3. Do the pitch balancing.
  4. Glue the second standoff in place. 

Other attachments:

Attachment #3 - Unglued stand off with wire in the groove, mirror freely suspended.

Attachment #4 - Glued stand off with wire in the groove, mirror freely suspended. Clearance between wire and magnet looks reasonable.

Attachment #5 - Barrel of optic (underside), mirror freely suspended. The wire seems to be in a reasonable orientation along the barrel, albeit not perfectly parallel.


Koji just pointed out that we should check that the unglued ruby standoff is in good contact with the barrel of the optic. Attachment #1 suggests that maybe this is not the case. If you zoom into Attachment #1, it is not clear if the standoff is sitting on the glue.

Attachment 1: IMG_2828.JPG  5.233 MB  | Hide | Hide all
IMG_2828.JPG
Attachment 2: IMG_2826.JPG  5.628 MB  | Hide | Hide all
IMG_2826.JPG
Attachment 3: IMG_2821.JPG  5.416 MB  | Hide | Hide all
IMG_2821.JPG
Attachment 4: IMG_2822.JPG  5.535 MB  | Hide | Hide all
IMG_2822.JPG
Attachment 5: IMG_2823.JPG  3.386 MB  | Hide | Hide all
IMG_2823.JPG
ELOG V3.1.3-